OpenStudy (anonymous):

find the equation of the line tangent to y=(3x/2+2)^2 at x=2.

OpenStudy (anonymous):

dy/dx=2*(3/2)*(3x/2+2) at x=2 dy/dx=15 at x=2 y=25 so y-25=15(x-2)

OpenStudy (anonymous):

what's the first step you did?

OpenStudy (anonymous):

find the derivative of y=(3x/2+2)^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

yes, thank you very much :) thanks to this site i will be able to get the subjects i dont get..i have my final on friday and i need to pass it

OpenStudy (anonymous):

you most welcome.......

OpenStudy (anonymous):

is the answer in this form (y=mx+b) this ... y=15x+20 ??

OpenStudy (anonymous):

no.. y-25=15x-30 y=15x-5

OpenStudy (anonymous):

why is it "-5" instead of "+20" ?

OpenStudy (anonymous):

how did you get y=15x+20..can u please show me..my calculation says its y=15x-5

OpenStudy (anonymous):

i solved for "b" ....after getting y'=15(which is what you got) i set equal to zero this..y'=3(3x/2+2) and got x=-4/3..i forgot how i got zero for "y" and then plugged in x and y and the slope and got b=20

OpenStudy (anonymous):

why u set equal to zero this..y'=3(3x/2+2) ??

OpenStudy (anonymous):

so i could get "x"